OpenStudy (anonymous):

The sides of a regular octagon are "s". Find a formula for the distance between the parallel sides of the octagon.

OpenStudy (amistre64):

Hmmm... an 8 sided polygon If the sides are equal to 1 then we can call it a unit octagon.... what are the interior angles of an octagon equal to?

OpenStudy (amistre64):

360/8 = 45 degrees 45/2 = 22.5 degrees tan(22.5) degrees = s/2x ; where x= half the distance across. 2x = s/tan(22.5) x= s*cot(22.5)/2 ; x is half the distance, so lets multiply both sides by "2" to get the full distance across d= s*cot(22.5). Thats the best I can come up with. the total distance across is equal to

OpenStudy (anonymous):

I don't have a problem with using trig to solve it. However this was in an algebra 2 text in a chapter using radicals in equations. the students haven't had trig. But thanks.